# Random Batch Sum-of-Gaussians Method for Molecular Dynamics of Born–Mayer–Huggins Systems

### v1

@@ Abstract
The Born–Mayer–Huggins (BMH) potential, which combines Coulomb interactions with dispersion and short-range exponential repulsion, is widely used for ionic materials such as molten salts. However, large-scale molecular dynamics simulations of BMH systems are often limited by computation, communication, and memory costs. We recently proposed the random batch sum-of-Gaussians (RBSOG) method, which accelerates Coulomb calculations by using a sum-of-Gaussians (SOG) decomposition to split the potential into short- and long-range parts and by applying importance sampling in Fourier space for the long-range part. In this work, we extend the RBSOG to BMH systems and incorporate a random batch list (RBL) scheme to further accelerate the short-range part, yielding a unified framework for efficient simulations with the BMH potential. The combination of the SOG decomposition and the RBL enables an efficient and scalable treatment of both long- and short-range interactions in BMH system, particularly the RBL well handles the medium-range exponential repulsion and dispersion by the random batch neighbor list. Error estimate is provided to show the theoretical convergence of the RBL force. We evaluate the framework on molten NaCl and mixed alkali halide with up to $5\times 10^{6}$ atoms on $2048$ CPU cores. Compared to the Ewald-based particle-particle particle-mesh method and the RBSOG-only method, our method achieves approximately $4\sim 10\times$ and $2\times$ speedups while using $1000$ cores, respectively, under the same level of structural and thermodynamic accuracy and with a reduced memory usage. These results demonstrate the attractive performance of our method in accuracy and scalability for MD simulations with long-range interactions.
@@
Born–Mayer–Huggins potential, molten salts, molecular dynamics simulations, sum-of-Gaussians, random batch list
@@ 1 Introduction
Here, $r_{ij}$ denotes the distance between ions $i$ and $j$; $q_{i}$ is the ionic charge; the exponential term models short-range Pauli repulsion with amplitude $A_{ij}$, range parameter $\rho$, and $\sigma_{ij}:=\sigma_{i}+\sigma_{j}$ being the addition of ionic radii; and the last two terms describe dipole–dipole and dipole–quadrupole dispersion with coefficients $C_{ij}$ and $D_{ij}$. Despite their broad use in molten-salt MD simulations [porter2022computational, galashev2023molecular, shore2000simulations], large-scale BMH simulations remain computationally costly. Several fundamental bottlenecks hinder large-scale BMH simulations. The long-range Coulomb term, typically handled by communication-intensive Ewald-type solvers with fast Fourier transforms (FFTs) [hockney2021computer, darden1993particle], introduces significant scalability limitations [ayala2021scalability]. Meanwhile, the exponential repulsion and dispersion terms exhibit slow algebraic decay, representing a medium-range interaction that requires a large real-space cutoff and converges slowly in Fourier space. Moreover, periodic fast multipole methods (FMMs) [yan2018flexibly, pei2023fast] further introduce complexity by splitting the periodic system into near‐field and far‐field regions; the near‐field interactions must be evaluated over the $27$ neighboring image cells, which is itself computationally costly. Recently proposed dual-space multilevel kernel-splitting (DMK) framework [jiang2025dual] and Ewald summation with prolates (ESP) method [liang2025accelerating] provide promising alternatives, but have not yet been adapted to BMH systems. These factors collectively limit both the accessible time scale and system size in BMH-based MD simulations.
@@
In recent years, stochastic algorithms have emerged and improved scalability by replacing global kernel sums with unbiased mini-batch estimators, and their accuracy has been rigorously established for interacting particle systems [jin2020random, jin2023ergodicity, ye2024error, cai2024convergence, lin2024hybrid]. For electrostatics, random batch Ewald (RBE) [jin2021random, liang2022superscalability, liang2022random] avoids global FFTs via importance sampling while keeping linear scaling. Since the construction of an SOG approximation has been extensively studied [greengard2018anisotropic, liang2020kernel, beylkin2005approximation], the random batch sum-of-Gaussians (RBSOG) method [liang2023random, chen2025random] has been proposed to replace the Ewald split with a sum-of-Gaussians (SOG) decomposition of $1/r$, removing the discontinuity of the Ewald split and further reducing batch size. However, the other terms of the BMH potential (exponential repulsion and dispersion) can remain the dominant cost because of their medium-range nature, creating a bottleneck that is not addressed by long-range solvers alone. The random batch list (RBL) scheme, developed to accelerate the short-range potentials such as the Lennard‑Jones (LJ) potential [liang2021random] and the embedded atom potential [zhang2025random], reduces the neighbor‑list cost by randomly sampling particles within a cutoff radius. The improved random batch Ewald (IRBE) approach [liangIRBE] combines RBE with the LJ-based RBL. Notably, the RBMD package [gao2026rbmd] implements the IRBE on GPU–CPU heterogeneous architectures, enabling efficient all-atom simulations of up to ten million particles on a single GPU. However, IRBE has not been applied to BMH systems and does not take advantage of the more efficient SOG decomposition. A linear-scaling method that treats all components of the BMH potential in a stochastic framework is of great significance for practical simulations.
@@
In order to overcome the challenge of handling medium-range interactions in BMH systems, we propose a unified framework, the improved random batch sum-of-Gaussians (IRBSOG) method, which seamlessly couples the RBSOG and RBL schemes. The key contributions of this work are threefold. First, we develop an integrated approach for the BMH potential in which the Coulomb kernel is split via a sum-of-Gaussians (SOG) approximation, with the long-range part efficiently evaluated via RBSOG in Fourier space using random batch importance sampling, while the short-range residual, together with the exponential repulsion and dispersion terms, is handled by a core-shell RBL scheme in real space. This synergistic combination specifically targets the multi-scale nature of the BMH potential, enabling scalable and accurate treatment of both interaction regimes. Second, recognizing that short-range interactions in BMH systems exhibit slower decay and require larger cutoffs compared to Lennard-Jones-type potentials, we systematically investigate and optimize the RBL implementation under these conditions. An adaptive, species-dependent cutoff strategy is introduced to account for size differences between alkali-metal cations and halogen anions, effectively reducing neighbor counts and memory usage. Third, we performed systematic MD simulations across a range of BMH systems—including molten NaCl and mixed alkali halides—to thoroughly validate the accuracy, scalability, and transferability of our approach. The proposed IRBSOG framework achieves $O(N)$ complexity with significantly reduced communication and memory overhead compared to both the particle-particle particle-mesh (PPPM) and the RBSOG-only methods. In large-scale benchmarks, it reproduces key structural, dynamical, and thermodynamic properties while delivering up to an order-of-magnitude speedup over PPPM and a roughly twofold improvement over RBSOG-only simulations for systems of up to $5\times 10^{6}$ atoms on $2048$ CPU cores.
@@ 2.3 Convergence analysis of IRBSOG-based simulations
Under the Debye–Hückel (DH) assumption, the variance of the long-range force approximation $\bm{F}_{\mathcal{F},i}^{*}$ satisfies
@@
We now analyze the convergence of IRBSOG-based MD. Let $\Delta t$ be the step size of MD time integration. Under the canonical (NVT) ensemble, a thermostat is generally used to sample correct ensemble distribution, such as the Langevin thermostat [frenkel2023understanding] and the Nosé–Hoover (NH) thermostat [hoover1985canonical]. We consider the Langevin thermostat for convenience in analysis. Let $(\bm{r}_{i},\bm{v}_{i})$ be the solution of the underdamped Langevin equations of motion,
@@
By Theorem 2.2, we have $\Lambda=O(1/P_{R}+1/P_{F})$. For long-time simulations, a uniform-in-time error estimate can be derived in a manner similar to [jin2022random] for systems with regular force fields. We note that Theorem 2.3 assumes the force to be bounded and Lipschitz continuous, whereas the Coulomb kernel has a singularity at the origin. Nevertheless, we expect that the error bound in Eq. (47) remains valid in practice. This is because the BMH potential includes both Lennard–Jones–type interactions and a strong short-range exponential repulsion, which prevent particles from approaching arbitrarily close, so that the singularity does not occur during the simulation. A fully rigorous justification of this argument remains difficult and highly nontrivial. For simulations in the microcanonical (NVE) ensemble, random-batch-type methods can be combined with a weakly coupled thermal bath to improve long-time energy stability [liang2024energy]. For further discussion of other thermostats and ensembles, we refer the reader to [jin2021random, liang2023random].
@@ 3.1 Accuracy comparison in molten NaCl systems
We further assess accuracy in fluid mechanics and thermodynamics by computing the viscosity and thermal conductivity. Viscosity is the internal resistance to flow and is obtained from the Green–Kubo relation in equilibrium [dunweg2008colloidal]:
@@
with $\bar{e}_{i}$ the per-atom energy. As shown in Fig. 2(c–d), IRBSOG with $P_{R}=30$ and $P_{F}=100$ yields viscosity and thermal conductivity for molten NaCl at 1100K1100~K and 1500K1500~K that are statistically indistinguishable from PPPM with $10^{-5}$ accuracy. This indicates that a modest batch size suffices for IRBSOG to produce accurate MD results. For reference, IRBE [liangIRBE] requires a Fourier-space batch size of $P_{F}=200$, twice that of the IRBSOG in simulating BMH systems, when simulating pure water and ionic liquid systems. The factor of two speedup comes from replacing the Ewald decomposition with the SOG decomposition.
@@ 3.3 Tests on molten alkali binary halide systems
Molten alkali binary halides, although composed of only three ionic species, show complex thermodynamic and transport behavior beyond that of simple molten salts or dilute electrolytes. Key properties—including melting point, viscosity, and electrical conductivity—can vary nonmonotonically with composition, reflecting strong ion–ion correlations and composition-dependent structural reorganization in the melt. Accurately capturing these effects is therefore a stringent test for interatomic models used at high temperature.
To assess the performance of the IRBSOG method, we conduct MD simulations of the LiCl–NaCl mixture over a wide range of compositions. The potential parameters are summarized in Table 3. For IRBSOG, we set $r_{\mathrm{core}}=5$$\text{\AA }$, $r_{\mathrm{shell}}=10$$\text{\AA }$, and use batch sizes $P_{R}=30$ and $P_{F}=100$. Figure 5 shows the RDFs for six representative compositions. IRBSOG reproduces composition-dependent shifts in peak positions and changes in coordination features across the first and second shells, as well as medium-range correlations. These results indicate that IRBSOG maintains accuracy and transferability for chemically mixed molten salts in which cross-species interactions play a central role.

### v2

# Random Batch Sum-of-Gaussians Method for Molecular Dynamics of Born-Mayer-Huggins Systems

@@ Abstract
The Born-Mayer-Huggins (BMH) potential, which combines Coulomb interactions with dispersion and short-range exponential repulsion, is widely used for ionic materials such as molten salts. However, large-scale molecular dynamics simulations of BMH systems are often limited by computation, communication, and memory costs. We recently proposed the random batch sum-of-Gaussians (RBSOG) method, which accelerates Coulomb calculations by using a sum-of-Gaussians (SOG) decomposition to split the potential into short- and long-range parts and by applying importance sampling in Fourier space for the long-range part. In this work, we extend the RBSOG to BMH systems and incorporate a random batch list (RBL) scheme to further accelerate the short-range part, yielding a unified framework for efficient simulations with the BMH potential. The combination of the SOG decomposition and the RBL enables an efficient and scalable treatment of both long- and short-range interactions in BMH system, particularly the RBL well handles the medium-range exponential repulsion and dispersion by the random batch neighbor list. Error estimate is provided to show the theoretical convergence of the RBL force. We evaluate the framework on molten NaCl and mixed alkali halide with up to $5\times 10^{6}$ atoms on $2048$ CPU cores. Compared to the Ewald-based particle-particle particle-mesh method and the RBSOG-only method, our method achieves approximately $4\sim 10\times$ and $2\times$ speedups while using $1000$ cores, respectively, under the same level of structural and thermodynamic accuracy and with a reduced memory usage. These results demonstrate the attractive performance of our method in accuracy and scalability for MD simulations with long-range interactions.
@@
Born-Mayer-Huggins potential, molten salts, molecular dynamics simulations, sum-of-Gaussians, random batch list
@@ 1 Introduction
Here, $r_{ij}$ denotes the distance between ions $i$ and $j$; $q_{i}$ is the ionic charge; the exponential term models short-range Pauli repulsion with amplitude $A_{ij}$, range parameter $\rho$, and $\sigma_{ij}:=\sigma_{i}+\sigma_{j}$ being the addition of ionic radii; and the last two terms describe dipole-dipole and dipole-quadrupole dispersion with coefficients $C_{ij}$ and $D_{ij}$. Despite their broad use in molten-salt MD simulations [porter2022computational, galashev2023molecular, shore2000simulations], large-scale BMH simulations remain computationally costly. Several fundamental bottlenecks hinder large-scale BMH simulations. The long-range Coulomb term, typically handled by communication-intensive Ewald-type solvers with fast Fourier transforms (FFTs) [hockney2021computer, darden1993particle], introduces significant scalability limitations [ayala2021scalability]. Meanwhile, the exponential repulsion and dispersion terms exhibit slow algebraic decay, representing a medium-range interaction that requires a large real-space cutoff and converges slowly in Fourier space. Moreover, periodic fast multipole methods (FMMs) [yan2018flexibly, pei2023fast] further introduce complexity by splitting the periodic system into near‐field and far‐field regions; the near‐field interactions must be evaluated over the $27$ neighboring image cells, which is itself computationally costly. Recently proposed dual-space multilevel kernel-splitting (DMK) framework [jiang2025dual] and Ewald summation with prolates (ESP) method [liang2025accelerating] provide promising alternatives, but have not yet been adapted to BMH systems. These factors collectively limit both the accessible time scale and system size in BMH-based MD simulations.
@@
In recent years, stochastic algorithms have emerged and improved scalability by replacing global kernel sums with unbiased mini-batch estimators, and their accuracy has been rigorously established for interacting particle systems [jin2020random, jin2023ergodicity, ye2024error, cai2024convergence, lin2024hybrid]. For electrostatics, random batch Ewald (RBE) [jin2021random, liang2022superscalability, liang2022random] avoids global FFTs via importance sampling while keeping linear scaling. Since the construction of an SOG approximation has been extensively studied [greengard2018anisotropic, liang2020kernel, beylkin2005approximation], the random batch sum-of-Gaussians (RBSOG) method [liang2023random, chen2025random] has been proposed to replace the Ewald split with a sum-of-Gaussians (SOG) decomposition of $1/r$, removing the discontinuity of the Ewald split and further reducing batch size. However, the other terms of the BMH potential (exponential repulsion and dispersion) can remain the dominant cost because of their medium-range nature, creating a bottleneck that is not addressed by long-range solvers alone. The random batch list (RBL) scheme, developed to accelerate the short-range potentials such as the Lennard‑Jones (LJ) potential [liang2021random] and the embedded atom potential [zhang2025random], reduces the neighbor‑list cost by randomly sampling particles within a cutoff radius. The improved random batch Ewald (IRBE) approach [liangIRBE] combines RBE with the LJ-based RBL. Notably, the RBMD package [gao2026rbmd] implements the IRBE on GPU-CPU heterogeneous architectures, enabling efficient all-atom simulations of up to ten million particles on a single GPU. However, IRBE has not been applied to BMH systems and does not take advantage of the more efficient SOG decomposition. A linear-scaling method that treats all components of the BMH potential in a stochastic framework is of great significance for practical simulations.
@@
In order to overcome the challenge of handling medium-range interactions in BMH systems, we propose a unified framework, the improved random batch sum-of-Gaussians (IRBSOG) method, which seamlessly couples the RBSOG and RBL schemes. The key contributions of this work are threefold. First, we develop an integrated approach for the BMH potential in which the Coulomb kernel is split via a sum-of-Gaussians (SOG) approximation, with the long-range part efficiently evaluated via RBSOG in Fourier space using random batch importance sampling, while the short-range residual, together with the exponential repulsion and dispersion terms, is handled by a core-shell RBL scheme in real space. This synergistic combination specifically targets the multi-scale nature of the BMH potential, enabling scalable and accurate treatment of both interaction regimes. Second, recognizing that short-range interactions in BMH systems exhibit slower decay and require larger cutoffs compared to Lennard-Jones-type potentials, we systematically investigate and optimize the RBL implementation under these conditions. An adaptive, species-dependent cutoff strategy is introduced to account for size differences between alkali-metal cations and halogen anions, effectively reducing neighbor counts and memory usage. Third, we performed systematic MD simulations across a range of BMH systems–including molten NaCl and mixed alkali halides–to thoroughly validate the accuracy, scalability, and transferability of our approach. The proposed IRBSOG framework achieves $O(N)$ complexity with significantly reduced communication and memory overhead compared to both the particle-particle particle-mesh (PPPM) and the RBSOG-only methods. In large-scale benchmarks, it reproduces key structural, dynamical, and thermodynamic properties while delivering up to an order-of-magnitude speedup over PPPM and a roughly twofold improvement over RBSOG-only simulations for systems of up to $5\times 10^{6}$ atoms on $2048$ CPU cores.
@@ 2.3 Convergence analysis of IRBSOG-based simulations
Under the Debye-Hückel (DH) assumption, the variance of the long-range force approximation $\bm{F}_{\mathcal{F},i}^{*}$ satisfies
@@
We now analyze the convergence of IRBSOG-based MD. Let $\Delta t$ be the step size of MD time integration. Under the canonical (NVT) ensemble, a thermostat is generally used to sample correct ensemble distribution, such as the Langevin thermostat [frenkel2023understanding] and the Nosé-Hoover (NH) thermostat [hoover1985canonical]. We consider the Langevin thermostat for convenience in analysis. Let $(\bm{r}_{i},\bm{v}_{i})$ be the solution of the underdamped Langevin equations of motion,
@@
By Theorem 2.2, we have $\Lambda=O(1/P_{R}+1/P_{F})$. For long-time simulations, a uniform-in-time error estimate can be derived in a manner similar to [jin2022random] for systems with regular force fields. We note that Theorem 2.3 assumes the force to be bounded and Lipschitz continuous, whereas the Coulomb kernel has a singularity at the origin. Nevertheless, we expect that the error bound in Eq. (47) remains valid in practice. This is because the BMH potential includes both Lennard-Jones-type interactions and a strong short-range exponential repulsion, which prevent particles from approaching arbitrarily close, so that the singularity does not occur during the simulation. A fully rigorous justification of this argument remains difficult and highly nontrivial. For simulations in the microcanonical (NVE) ensemble, random-batch-type methods can be combined with a weakly coupled thermal bath to improve long-time energy stability [liang2024energy]. For further discussion of other thermostats and ensembles, we refer the reader to [jin2021random, liang2023random].
@@ 3.1 Accuracy comparison in molten NaCl systems
We further assess accuracy in fluid mechanics and thermodynamics by computing the viscosity and thermal conductivity. Viscosity is the internal resistance to flow and is obtained from the Green-Kubo relation in equilibrium [dunweg2008colloidal]:
@@
with $\bar{e}_{i}$ the per-atom energy. As shown in Fig. 2(c-d), IRBSOG with $P_{R}=30$ and $P_{F}=100$ yields viscosity and thermal conductivity for molten NaCl at 1100K1100~K and 1500K1500~K that are statistically indistinguishable from PPPM with $10^{-5}$ accuracy. This indicates that a modest batch size suffices for IRBSOG to produce accurate MD results. For reference, IRBE [liangIRBE] requires a Fourier-space batch size of $P_{F}=200$, twice that of the IRBSOG in simulating BMH systems, when simulating pure water and ionic liquid systems. The factor of two speedup comes from replacing the Ewald decomposition with the SOG decomposition.
@@ 3.3 Tests on molten alkali binary halide systems
Molten alkali binary halides, although composed of only three ionic species, show complex thermodynamic and transport behavior beyond that of simple molten salts or dilute electrolytes. Key properties–including melting point, viscosity, and electrical conductivity–can vary nonmonotonically with composition, reflecting strong ion-ion correlations and composition-dependent structural reorganization in the melt. Accurately capturing these effects is therefore a stringent test for interatomic models used at high temperature.
To assess the performance of the IRBSOG method, we conduct MD simulations of the LiCl-NaCl mixture over a wide range of compositions. The potential parameters are summarized in Table 3. For IRBSOG, we set $r_{\mathrm{core}}=5$$\text{\AA }$, $r_{\mathrm{shell}}=10$$\text{\AA }$, and use batch sizes $P_{R}=30$ and $P_{F}=100$. Figure 5 shows the RDFs for six representative compositions. IRBSOG reproduces composition-dependent shifts in peak positions and changes in coordination features across the first and second shells, as well as medium-range correlations. These results indicate that IRBSOG maintains accuracy and transferability for chemically mixed molten salts in which cross-species interactions play a central role.